A straight line L is parpendicular to the line 5x-y=1. The area of the triangle formed by the line L and the coordinate axes is 5. Find the equation of the line L.

Grade:10

1 Answers

Ritesh Khatri
6 years ago
line parpendicular to given line is x + 5y = a. 
this can be written as x/a + y/(a/5) = 1
area  = ½ * base * height
 = (1/2)* a * (a/5) = 5
a^2 = 50
a = +(50^(1/2) , – (50^(1/2)
therefore equation of line is x + 5y = +(50^(1/2) and x + 5y = – (50^(1/2)